The time it takes for seed of a certain variety to germinate is normally distributed with a mean of 120 hours and a standard deviation of 20 hours. What is the probability that a randomly selected seed takes more than 130 hours to germinate?

A.) 0.49

B.) 0.6915

C.) 0.3085

D.) .025

E.) 0.508

F.) 0

G.) 0.5
Answer #1

Solution :

Given that ,

P(x > 130) = 1 - P(x < 130)

= 1 - P[(x - ) / < (130 - 120) / 20)

= 1 - P(z < 0.5)

= 1 - 0.6915

= 0.3085

Probability = 0.3085

option C.) is correct